OpenStudy (anonymous):

A 70 Kg person is parachuting and experiencing a downward acceleration of 3m/s. The mass of the parachute is 5kg. The upward force on the open parachute by the air is take g=10m/sec^2. a) 525N b)650 c)760 d)143

OpenStudy (irishboy123):

F = ma here you need F= net force; and m = 70+5 and they are both accelerating.....

OpenStudy (anonymous):

Could you elaborate the given clue to solution that would be helpful

OpenStudy (michele_laino):

on the system: person + parachute, are acting two forces: the total weight force whose magnitude is \(F=(70+5) \cdot 10=...\) and the resistance force \(R\) Now, using the second law of Newton, we can write this: \[\huge F - R = \left( {M + m} \right)a\] where \(a=3\), and \(M+m=70+5\), namely it is the mass of the system: person + parachute Please, solve such equation for \(R\)

OpenStudy (brrandyn):

Brilliant, @Michele_Laino! Thoroughly explained.

OpenStudy (michele_laino):

thanks! :) @Brrandyn

OpenStudy (anonymous):

Awesome explanation! @Michele_Laino

OpenStudy (michele_laino):

thanks! :) @Robert136

OpenStudy (anonymous):

k